OpenStudy (anonymous):

how do i write a function for this? the area A(n) of a square when you know the length (n) of a side

OpenStudy (anonymous):

A(n) = n*n = n^2

OpenStudy (turingtest):

A=LxW for a SQUARE L=W so A=L^2 if L=n A(n)=n^2